# Satellite image classification with neural quantum kernels

### v1

@@ III.2 Image pre-processing
The encoder produces 64-sized vectors as output from the bottleneck dense layer. Depending on the size of the parameterized circuit, these features can be further reduced through dimensionality reduction techniques. The pipeline selected for this problem is displayed in Figure 2. Once the latent features have been obtained, $z$-score normalization is applied to them prior to the application of dimensionality reduction through PCA. The number of components $p$ for PCA is selected as a hyperparameter of the model that depends on the specific requirements of the circuit’s encoding. Before embedding the features into the quantum circuit, their values are scaled to fit within the range of $[-1,1]$. Finally, the circuit yields the prediction $\hat{y}$, that is either positive class +1 or negative -1.
Figure 4: Detailed NQK construction process. The diagram starts with the training of a single-qubit QNN, from which we obtain the optimal parameters $\vec{\theta}^{*}$. Using these parameters, a $1$-to-$n$ NQK can be constructed. The upper part of the figure illustrates the specific case of a $1$-to-$2$ NQK construction. From the trained single-qubit QNN, we can also scale up to construct a 2-qubit QNN, as shown in the lower part of the figure. By iteratively applying this strategy, an $n$-qubit QNN can be constructed and used as the embedding for an $n$-to-$n$ NQK.
@@ IV Methodology
Classical pre-processing serves as a vital step in handling highly intricate real-world datasets. Within our pre-processing framework detailed in Section III.2, PCA stands out as a key technique which inherently introduces nonlinearities to the model. This way, the choice of the number of principal components $p$ becomes a crucial hyperparameter of the model.
In our study, we focus on $p=2$ and $p=3$ features since we are using qubits for the quantum part. Additionally, despite retaining more data information, considering a larger number of principal components does not guarantee improved performance, as our findings will illustrate.
@@ IV.1 Data re-uploading QNN
This structure alternates parametrized gates $U(\vec{\theta}_{l})$ with encoding gates $U(\vec{x})$. Here, $U$ represents a generic $\mathrm{SU}(2)$ unitary, parameterized by three angles. The number of free parameters is $3\times L$, where $L$ denotes the number of layers, contained within the vector $\vec{\theta}=\{\vec{\theta}_{1},...,\vec{\theta}_{L}\}$. Since we are considering data points $\vec{x}$ with $p=2$ and $p=3$ features, we only need a single encoding unitary. For data points with $p=2$ features, we set the third angle of the encoding gate to 0. This architecture yields highly expressive models, as demonstrated by the analyzed Fourier frequencies of the functions generated by the model Schuld et al. (2021); Casas and Cervera-Lierta (2023). Remarkably, even with a single qubit, one can perform complex classification tasks Tapia et al. (2023); Easom-McCaldin et al. (2024).
@@
[…]
@@
Figure 5: Numerical results for the $1$-to-$n$ NQKs approach. In $(a)$ and $(b)$, we consider the $1$-to-$2$ NQK, while in $(c)$ we examine the $1$-to-$3$ NQK. The results are displayed for $p=2$ features. The training and test accuracies are drawn from the 10-fold process using the described “$1$-to-$n$ set” of 2000 samples. In the box plots, the upper and lower parts of the boxes represent the 25th and 75th percentiles of the accuracies obtained by evaluating the accuracy of the classifiers using 10-fold. The horizontal line inside the box indicates the median, while the whiskers extend from the box to the most distant evaluation score that lies within 1.5 times the interquartile range (IQR) from the box. Score values beyond the whiskers are considered outliers.
@@ V Numerical results and analysis
Figure 6: Numerical results for the $n$-to-$n$ architecture, tested using datasets with 500 training points and 200 test points. Experiments were conducted using 2 and 3 features, with results averaged over five independent trials. The “scale QNN set” was utilized for these experiments, as described in the main text. The figure displays the average training and test accuracies for both the QNN and the corresponding $n$-to-$n$ NQK.
@@ V.1 $1$-to-$n$ NQK
We examine both $p=2$ and $p=3$ features but primarily focused on $p=2$, as using three features surprisingly performs less effectively for this classification problem. The results for $p=3$ are presented in Appendix C. For the number of qubits in the EQK, we considered $n=2$ and $n=3$, corresponding to $1$-to-$2$ and $1$-to-$3$ NQKs, respectively. For this classification task, increasing the number of qubits does not enhance performance. Notably, we achieved high classification accuracies even with a small number of qubits.
@@ V.3 Classical benchmark
Model $n$ components Train Test
@@
Until now, our investigation has revolved around a hybrid architecture combining classical preprocessing with two distinct QML methodologies. Presently, we delve into classical approaches as substitutes for the quantum components to benchmark our models. In order to accomplish this, we use the preprocessing framework outlined in Section III.2 and then perform the same 10-fold approximation for training and evaluation with classical machine learning algorithms.
@@
The selected models for benchmarking are Support Vector Machine classifier (SVC) and Random Forest classifier. As shown in Table 1, three runs are performed for each model. In order to ensure a fair comparison with the EQK models, the number of features is fixed to $p=2$ and $p=3$ in the first two rows for each model. A hyperparameter optimization with randomized search is then performed to yield nearly optimal test results. In the last row for each model, also $p$ is treated as an hyperparameter in the randomized search to determine if it would yield better results than using $p\in\{2,3\}$. The results of our analysis indicate that the optimal value of $p$ for the SVC is $p=45$, while for the Random Forest it is $p=10$. The values for the remaining hyperparameters for each model can be found in Appendix D.
VI Conclusions
This work addresses a complex binary classification task on a real dataset, focusing on the classification of satellite images to determine whether they contain solar panels or not. Classical preprocessing is conducted to reduce the data dimensionality, employing a cross-task transfer learning implementation, which demonstrates notable effectiveness for this purpose. Once the data dimensionality is reduced to enable processing with quantum models, we utilize two different NQK approaches achieving maximal accuracies near 90%.
We initially explored the $1$-to-$n$ NQK approach. In this model, we observed that constructing an EQK from the training of a single-qubit QNN demonstrates robustness, yielding nearly identical results whether the QNN is optimally or sub-optimally trained. Additionally, we investigated $n$-to-$n$ NQKs, which involve building a scalable $n$-qubit QNN and constructing the corresponding quantum kernel. During the scaling of the QNN, we found that the highest accuracies were achieved by increasing the number of qubits. We also observed robustness in the results when constructing the kernel. With both methods, we successfully match the performance of classical methods, which have also been optimized for the number of features. While the results presented focus on a limited number of qubits, suggesting that the quantum models could be considered quantum-inspired, we demonstrate that, for this classification task, the accuracy values achieved using various quantum and classical methods converge to similar levels.
This work aims to solve a real classification problem with industrial applications using quantum methodology. For this task, we have observed that with a low number of qubits, we can achieve accuracies comparable to the best classical models. However, it would be of great interest to explore tasks that require further scaling of these approaches in terms of the number of qubits. For future considerations, we could assess the performance of this approach on datasets with greater class imbalance and consider implementing it in on noisy quantum simulation backends or actual quantum hardware.
VII Acknowledgements
@@
This work was supported by the Spanish Ministry of Science and Innovation under the Recovery, Transformation and Resilience Plan (CUCO, MIG-20211005). PR and MS acknowledge support from HORIZON-CL4- 2022-QUANTUM01-SGA project 101113946 OpenSuperQPlus100 of the EU Flagship on Quantum Technologies, the Spanish Ramón y Cajal Grant RYC-2020-030503-I, project Grant No. PID2021-125823NA-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” and “ERDF Invest in your Future”, and from the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government. This project has also received support from the Spanish Ministry of Economic Affairs and Digital Transformation through the QUANTUM ENIA project call - Quantum Spain, and by the EU through the Recovery, Transformation and Resilience Plan - NextGenerationEU. We acknowledge funding from Basque Government through Grant No. IT1470-22 and through the ELKARTEK program, project ”KUBIT - Kuantikaren Berrikuntzarako Ikasketa Teknologikoa” (KK-2024/00105).
@@ Appendix B Dataset splits
Appendix C Numerical experiments
Presented here are additional numerical findings not featured in the main text. In Figure C.1, we explore all the NQK construction scenarios, encompassing $p=2$ and $p=3$ features, optimal and sub-optimal QNN training methods, and $n=2$ and $n=3$ qubits. As discussed in the main text, employing $p=2$ features yields superior performance for this architecture and classification task. We can also observe no improvement when transitioning from $n=2$ to $n=3$ qubits in this context.
@@ Appendix D Hyperparameters for the classical benchmark
Model n_components kernel C gamma
@@
Model n_components n_estimators max_depth max_features

### v2

@@ III.2 Feature engineering
The encoder produces 64-sized vectors as output from the bottleneck dense layer. Depending on the size of the parameterized circuit, these features can be further reduced through dimensionality reduction techniques. Once the latent features have been obtained, $z$-score normalization is applied to them prior to the application of dimensionality reduction through PCA. The number of components $p$ for PCA is selected as a hyperparameter of the model that depends on the specific requirements of the circuit’s encoding. Before embedding the features into the quantum circuit, their values are scaled to fit within the range of $[-1,1]$. Finally, the circuit yields the prediction $\hat{y}$, that is either +1 (positive class) or -1 (negative class).
It is important to mention that a large variety of alternative feature extractions techniques might also have been considered, such as: a) classical computer vision features extraction techniques applied at image level (e.g. SIFT, HOG, etc.), b) object-based image analysis techniques based on the sequential application of unsupervised segmentation methods (such as Quickshift, Felzenszwalb, SNIC, SLIC, etc.) followed by extraction of features related to shapes, colors, and other useful image patterns (e.g. GLCM, Hough transform, etc) , c) the application of other deep learning techniques adopting a neural network for extracting features in a more end-to-end, task oriented fashion (e.g., VGGNet, ResNet, EfficientNet, Vision Transformers, etc.). While some of them may also result to be effective for the task at hand, a comparison of alternative feature extraction methods is not implemented in this work that instead focuses on the application of the proposed hybrid QML algorithms, and their variants, to the selected complex real dataset use case.
@@ IV Methodology
Classical pre-processing serves as a vital step in handling highly intricate real-world datasets. Within our pre-processing framework detailed in Section III, PCA stands out as a key technique which inherently introduces nonlinearities to the model. This way, the choice of the number of principal components $p$ becomes a crucial hyperparameter of the model.
In our study, we focus on $p=2$ and $p=3$ features since we are using qubits for the quantum part. Additionally, despite retaining more data information, considering a larger number of principal components does not guarantee improved performance, as shown in the Appendix.
@@
Figure 4: Detailed NQK construction process. The diagram begins with the training of a single-qubit QNN, from which the optimal parameters $\vec{\theta}^{*}$ are obtained. Using these parameters, two types of NQKs can be constructed. The first is the $1$-to-$n$ NQK, which is shown in the upper part of the figure, illustrating the specific case where $n=2$. The second is the $n$-to-$n$ NQK, where we iteratively add qubits to construct an $n$-qubit QNN. The first step in constructing the 2-qubit QNN is shown in the lower-left part of the figure. This $n$-qubit QNN is then used directly as the embedding to create the $n$-to-$n$ NQK, as shown in the lower-right part.
@@ IV.1 Data re-uploading QNN
This structure alternates parametrized gates $U(\vec{\theta}_{l})$ with encoding gates $U(\vec{x})$. Here, $U$ represents a generic $\mathrm{SU}(2)$ unitary, parameterized by three angles. The number of free parameters is $3\times L$, where $L$ denotes the number of layers, contained within the vector $\vec{\theta}=\{\vec{\theta}_{1},...,\vec{\theta}_{L}\}$. With a single unitary we can encode up to $p=3$ features. For data points with $p=2$ features, we set the third angle of the encoding gate to 0. This architecture yields highly expressive models, as demonstrated by the analyzed Fourier frequencies of the functions generated by the model Schuld et al. (2021); Casas and Cervera-Lierta (2023). Remarkably, even with a single qubit, one can perform complex classification tasks Tapia et al. (2023); Easom-McCaldin et al. (2024).
@@
$\hat{y}[\vec{x}_{t}]=\mathrm{sign}\Bigl{(}\mathrm{tr}(|0\rangle\langle 0|\;\rho_{\vec{\theta}^{*}}(\vec{x}_{t}))-1/2\Bigr{)}\in\{-1,+1\},$ (5)
@@ IV.2 Neural quantum kernels
Figure 4 illustrates the entire process of constructing NQKs for both the $1$-to-$n$ and $n$-to-$n$ cases. The process begins with the training of a single-qubit neural network (top left), where we obtain the optimal parameters $\vec{\theta}^{*}$. Once we have these parameters, two construction approaches are possible. The $1$-to-$n$ approach, exemplified for the case $n=2$ in the top right, involves replicating the architecture of the trained single-qubit QNN and adding CNOT gates between the two qubits in each layer. This results in a 2-qubit EQK, referred to as a 1-to-2 NQK. The $n$-to-$n$ approach, described at the bottom of the figure, requires constructing an $n$-qubit QNN through a scalable iterative process explained earlier in this section. Once the QNN is built, it is used as an embedding to generate an $n$-qubit EQK, as shown in the bottom right.
@@
Figure 5: Numerical results for the $1$-to-$n$ NQKs approach. $1q$, $2q$, and $3q$ refer to one-qubit, two-qubit, and three-qubit constructions, respectively. In $(a)$ and $(b)$, we consider the $1$-to-$2$ NQK, while in $(c)$ we examine the $1$-to-$3$ NQK. The results are displayed for $p=2$ features. The training and test accuracies are drawn from the 10-fold process using the described “$1$-to-$n$ set” of 2000 samples. In the box plots, the upper and lower parts of the boxes represent the 25th and 75th percentiles of the accuracies obtained by evaluating the accuracy of the classifiers using 10-fold. The horizontal line inside the box indicates the median, while the whiskers extend from the box to the most distant evaluation score that lies within 1.5 times the interquartile range (IQR) from the box. Score values beyond the whiskers are considered outliers.
@@ V.1 $1$-to-$n$ NQK
We examine both $p=2$ and $p=3$ features but primarily focused on $p=2$, as using three features performs less effectively with this model. The results for $p=3$ are presented in Appendix C. For the number of qubits in the EQK, we considered $n=2$ and $n=3$, corresponding to $1$-to-$2$ and $1$-to-$3$ NQKs, respectively. For this classification task, increasing the number of qubits does not enhance performance. Notably, we achieved high classification accuracies even with a small number of qubits.
@@
Figure 6: Numerical results for the $n$-to-$n$ architecture, tested using datasets with 500 training points and 200 test points. The experiments show the mean results from five independent trials, with 2 features displayed in panels (a) and (b), and 3 features in panels (c) and (d). The “scale QNN set”, which is described in the main text, was utilized for these experiments. The figure displays the average training and test accuracies for both the QNN and the corresponding $n$-to-$n$ NQK.
@@ V.3 Classical benchmark
Model # features ($p$) Training accuracy Test accuracy
@@
Until now, our investigation has revolved around a hybrid architecture combining classical preprocessing with two distinct QML methodologies. Presently, we delve into classical approaches as substitutes for the quantum components to benchmark our models. In order to accomplish this, we use the preprocessing framework outlined in Section III and then perform the same 10-fold approximation for training and evaluation with classical machine learning algorithms.
@@
The selected models for benchmarking are Support Vector Machine classifier (SVC) and Random Forest classifier. As shown in Table 1, three runs are performed for each model. In order to ensure a fair comparison with the NQK models, the number of features is fixed to $p=2$ and $p=3$ in the first two rows for each model. A hyperparameter optimization with randomized search is then performed to yield nearly optimal test results. In the last row for each model, also $p$ is treated as an hyperparameter in the randomized search to determine if it would yield better results than using $p\in\{2,3\}$. The results of our analysis indicate that the optimal value of $p$ for the SVC is $p=45$, while for the Random Forest it is $p=10$. The values for the remaining hyperparameters for each model can be found in Appendix D.
VI Discussion and limitations
This work highlights some conclusions from previous studies while also paving the way for new insights. As observed, the feature selection and classical preprocessing pipeline significantly influences the overall performance of the model, a finding consistent with prior literature Zollner et al. (2024); Otgonbaatar and Datcu (2022). Our results further emphasize that, with effective classical preprocessing, quantum models can be competitive in addressing real-world classification problems. Notably, these results are achievable with a limited number of qubits, making such models potentially quantum-inspired, as they can be directly simulated on classical computers, which opens the door to the development of new classical models.
However, as shown in our classical benchmark models, the findings also underline that classical models remain highly competitive in terms of performance. This suggests that the benefits of using quantum or quantum-inspired models may not lie solely in accuracy but in other metrics such as a reduced number of parameters or improved generalization capabilities, which merit further investigation.
Despite these promising insights, this work has limitations that should be addressed in future research. First, the model is restricted to a very small number of features. As observed, the straightforward method of adding features beyond proves ineffective, highlighting the necessity of exploring alternative approaches. A promising direction could involve leveraging qudits, which offer more degrees of freedom and could enable the encoding of additional features through single-qudit gates Wach et al. (2023); Roca-Jerat et al. (2024).
Another significant limitation is the inherent hardware noise in quantum processing units. While simulations have been used here as an initial proof of concept to showcase the competitiveness of quantum models, future studies should evaluate the effects of noise and the robustness of the models under realistic hardware conditions. Furthermore, if the addition of more features requires additional quantum gates, it becomes even more crucial to study the effects of hardware noise, as more gates naturally introduce more noise. Additionally, if more features necessitate the use of more qubits, the current limitations of quantum hardware size must be carefully considered.
VII Conclusions and future work
This work demonstrates that quantum models, when combined with effective classical preprocessing, can achieve competitive performance on a real-world binary classification task. Our results show that quantum models based on NQKs achieve accuracies close to 90%, rivaling well-optimized classical methods. We explored two NQK approaches: the $1$-to-$n$ method, which uses a single-qubit QNN and proves particularly robust, and the $n$-to-$n$ method, which improves accuracy as the number of qubits increases, addressing the scalability challenges commonly discussed in quantum machine learning.
Although quantum models show promise, classical models remain highly competitive in terms of accuracy, suggesting that the real advantages of quantum or quantum-inspired methods may lie in metrics beyond raw performance. Future work should focus on scaling these models to more qubits, addressing noise in quantum hardware, and exploring the potential of qudits for handling larger feature sets. This research opens the door for further exploration of quantum techniques in real-world industrial applications.
VIII Acknowledgements
@@
The authors would like to thank A. Benítez-Buenache for his helpful comments during the writing of the manuscript. This work was supported by the Spanish Ministry of Science and Innovation under the Recovery, Transformation and Resilience Plan (CUCO, MIG-20211005). PR and MS acknowledge support from HORIZON-CL4- 2022-QUANTUM01-SGA project 101113946 OpenSuperQPlus100 of the EU Flagship on Quantum Technologies, the Spanish Ramón y Cajal Grant RYC-2020-030503-I, project Grant No. PID2021-125823NA-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” and “ERDF Invest in your Future”, and from the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government. This project has also received support from the Spanish Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain, EU through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026. We acknowledge funding from Basque Government through Grant No. IT1470-22 and through the ELKARTEK program, project ”KUBIT - Kuantikaren Berrikuntzarako Ikasketa Teknologikoa” (KK-2024/00105).
@@ Appendix B Dataset splits
Appendix C Additional numerical experiments
C.1 $1$-to-$n$
@@
Presented here are additional numerical findings not featured in the main text. In Figure C.1, we explore all the $1$-to-$n$ NQK construction scenarios, encompassing $p=2$ and $p=3$ features, optimal and sub-optimal QNN training methods, and $n=2$ and $n=3$ qubits. As discussed in the main text, employing $p=2$ features yields superior performance for this architecture and classification task. We can also observe no improvement when transitioning from $n=2$ to $n=3$ qubits in this context.
C.2 Different number of features
As argued in the main text, when we use single-qubit unitaries to encode features, we are limited to encoding only 3 features with a single encoding unitary. For cases where we have $p>3$ features, as proposed in Ref. Pérez-Salinas et al. (2020), we require a number of encoding unitaries equal to $\lceil p/3\rceil$. For example, if we consider $p=6$ with a data sample $\bm{x}=(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})$, instead of using a single encoding unitary, we would need two—namely, $U(x_{1},x_{2},x_{3})\cdot U(x_{4},x_{5},x_{6})$.
However, a single qubit has a limited capacity for processing information, meaning that even with this approach of adding more features, the qubit’s processing ability becomes a bottleneck. Including additional features merely introduces more information than the qubit can handle effectively and increases the circuit depth without necessarily improving model performance. This is evident in Figure C.2, which is intended to clarify why we focus only on $p=2$ and $p=3$. The figure plots the mean training and test accuracies over the 5 dataset samples considered in Figure 6 of the $n$-qubit QNN.
As we can see, the best performance occurs with $p=3$, followed by $p=2$, where we are still able to use a single encoding unitary. Beyond this, adding more features degrades the results. For training accuracy, performance consistently drops as the number of features increases to $p=4,5,6$, whereas for test accuracy, $p=6$ performs slightly better than $p=5$.
Figure C.2: Mean training (left) and test (right) accuracies across 5 dataset samples for different number of features $p$ using the $n$-qubit QNN architecture described in the main text.
C.3 Different dimensionality reduction techniques
In the main text, we used PCA as the feature reduction technique. Here, we present overlaid results for $p=2$ and $p=3$ features obtained with PCA, along with two additional techniques from scikit-learn Python library: Independent Component Analysis (ICA) and Truncated Singular Value Decomposition (SVD).
Figure C.3 shows the results for the $n$-qubit QNN, presenting the average training and test accuracies across the same five dataset samples used in Figure 6. As observed, PCA outperforms the other two techniques in most cases, except for the test accuracies with $p=2$ features, where the results across all three techniques are very similar.
Figure C.3: Mean training and test accuracies across five dataset samples from the “$n$-to-$n$” set (described in the main text), for $p=2$ and $p=3$ features. The results are presented for three dimensionality reduction techniques: Independent Component Analysis (ICA), Truncated Singular Value Decomposition (SVD), and Principal Component Analysis (PCA), utilizing the $n$-qubit QNN architecture outlined in the main text.
Figure C.4 shows the results for the $1$-to-$2$ NQK approach with optimal training, using the two additional dimensionality reduction techniques, ICA and SVD. The dataset splits are the same as those used to generate Figure 5. As we can see, and when comparing with Figure C.1, the results using the three dimensionality reduction techniques are quite similar for both $p=2$ and $p=3$ features.
Figure C.4: Numerical results for the $1$-to-$2$ NQKs approach with optimal training, incorporating two additional dimensionality reduction techniques: Independent Component Analysis (ICA) and Truncated Singular Value Decomposition (SVD). The results are presented for $p=2$ and $p=3$ features. Training and test accuracies are obtained from a 10-fold cross-validation process using the described “$1$-to-$n$” dataset, consisting of 2000 samples. The box plots display the distribution of accuracies, where the upper and lower edges of the boxes represent the 25th and 75th percentiles, respectively. The horizontal line inside the box marks the median accuracy. The whiskers extend from the box to the most extreme evaluation score within 1.5 times the interquartile range (IQR). Any scores beyond the whiskers are considered outliers.
With these additional experiments, we aim to demonstrate that, while the choice of dimensionality reduction method does affect the results, selecting PCA as used in the main experiments is both reasonable and representative for conducting our study.
@@ Appendix D Hyperparameters for the classical benchmark
Model # features ($p$) kernel C gamma
@@
Model # features ($p$) n_estimators max_depth max_features